Sulphuric acid has a molecular mass of 98 and a density of 1.94 g/mL. You are given a 10% v/v solution of sulphuric acid. What is its molar concentration? Give your answer to 2 decimal places.
1
Expert's answer
2017-11-22T13:40:06-0500
If we have 1L of solution, then we have 100 ml (0.1L = 10% from 1L) of sulphuric acid in it.
Find the mass of 100 ml of sulphuric acid:
m = 100 × 1.94 = 194 (g)
Find the amount of sulphuric acid in 1L of solution:
ν = 194/98 = 1.98 (mole)
Answer
The molar concentration is 1.98 mole/L.